A student studying for a vocabulary test knows the meanings of 10 words from a list of 24 words. If the test contains 10 words from the study list, what is the probability that at least 8 of the words on the test are words that the student knows? (Round your answer to three decimal places.)
Answer #1

Binomial Distribution

n = 10

p = 10/24 = 0.4167

q = 1 - p = 0.5833

P(X\geq8) = P(X = 8) + P(X = 9) + P(X = 10)

x 0.58332 x 0.41678 = 45 x 0.34023889 x 0.0009091 P(X = 8) = 1 0.0139

P(X = 9) = 10 9 x0.58331 x 0.4167° = 10 x 0.5833 x 0.0003788 = 0.0022

P(X = 10) = 1 x 0.58330 x 0.416710 = 1x1x0.0001578 = 0.0001578

So,

P(X\geq8) = 0.0139 + 0.0022 + 0.0001578 = 0.016

So,

Answer is:

0.016